# RobustSpring: Benchmarking Robustness to Image Corruptions for Optical Flow, Scene Flow and Stereo

### v1

@@ Appendix A Repository Links and Checkpoints
Table A1: Repositories and checkpoints used for evaluating methods in RobustSpring. *The Mixed checkpoint MS-RAFT+ is pretrained on Chairs and Things and then fine-tuned on a mix of Sintel, Viper, KITTI 2015, Things, and HD1k.
Appendix B RobustSpring Image Corruptions
Below, we provide supplementary information on the image corruptions for the RobustSpring dataset. Besides visualizing further benchmark samples, we also give details on their implementation with a focus on RobustSpring-specific consistencies.
B.1 Additional Corruption Benchmark Samples
To complement the benchmark samples in Fig. 3, we show benchmark results on additional corruptions in Fig. A1.
@@ B.1 Additional Corruption Benchmark Samples
Figure A1: RobustSpring example frames, complementing Fig. 3. The first row shows clean and corrupted images. The second row shows the left and right disparity maps predicted with LEA Stereo [13]. The third row shows the target disparities for forward left, backward left, forward right, and backward right directions from M-FUSE [42]. The fourth row shows optical flow estimates for forward left, backward left, forward right, and backward right from RAFT [69]. All disparities and flows are computed on the corrupted dataset.
B.2 Corruption Implementation
@@
[…]
@@
[…]
@@
[…]
@@
Snow and Rain. The implementation for snow and rain is based on [61], with methodological and performance improvements. On the methodological side, we replaced additive blending with order-independent alpha blending (Meshkin’s Method) [41] and included global illumination [19] in the color rendering. Also, we expanded the monocular two-step motion simulation to multi-step stereo images. On the performance side, we introduce an efficient parallel particle initialization and improve the parallel processing performance.
Fog. Fog is modelled using the Koschmieder model from [73] as
@@
Figure A2: Accuracy vs. robustness of optical flow methods, measured as EPE and median $R^{c}_{\text{EPE}}$. Small values indicate accurate and robust methods. Fig. 5 shows the average $R^{c}_{\text{EPE}}$.
@@
Figure A3: Additional results on accumulated corruption robustness $R^{c}_{\text{EPE}}$ for optical flow models over corruption classes color, blur, quality, and weather. More results are in Fig. 4.
Appendix C Additional Experimental Results
@@
C.1 Corruption Robustness by Corruption Group
Figure A3 shows the corruption robustness $R^{c}_{\text{EPE}}$ for each optical flow method across the remaining four corruption groups in addition to Fig. 4 in the main paper. It underlines the varying degrees of robustness of the evaluated methods against specific types of corruption.
C.2 Accuracy vs. Median Corruption Robustness
In Fig. A2 we show the accuracy-robustness evaluation with the Median corruption robustness, to complement Fig. 5 which uses the average corruption robustness. Even though the robustness ranking of methods varies between average and media corruption robustness, cf. Sec. 4.2 for a discussion on ranking differences, the general trend that corruption robustness and accuracy are weakly correlated remains. However, there is still no clear winner, and an accuracy-robustness tradeoff persists among particularly accurate or robust methods.
C.3 Schulze Pairwise Comparison Matrix
The Schulze method is a ranking algorithm used to determine the most preferred candidate based on pairwise comparisons. We include a pairwise comparison matrix in Table A2 for our ranking. The table shows how often the method in row i is better than the method in column j, based on the number of corruptions where method i achieves a lower error than method j. The ranking process consists of the following steps:
@@ C.3 Schulze Pairwise Comparison Matrix
GMFlow 0 9 7 16 10 16 14 12
GMA 11 0 9 19 12 17 14 14
MS-RAFT+ 13 11 0 19 12 18 15 14
FlowFormer 4 1 1 0 8 9 13 8
SPyNet 10 8 8 12 0 12 14 5
RAFT 4 3 2 11 8 0 14 8
PWCNet 6 6 5 7 6 6 0 4
FlowNet2 8 6 6 12 15 12 16 0
Table A2: Pairwise comparison matrix for the Schulze method.
C.4 Corruption Robustness on Snow
@@
Figure A4: Stability of corruption robustness $R^{c}_{\text{EPE}}$ for snow corruption. Analogous rain results in Fig. 6.
@@
Finally, we complement the evaluation of our corruption robustness metric in the presence of rain in Fig. 6 with the corresponding evaluation in the presence of snow in Fig. A4. The results with rain translate to snow with the following minor differences: Because snow has less motion blur than rain, it covers fewer pixels (60% of all pixels vs. 90% for rain). For snow, the score drops a bit more than for rain when object pixels are excluded ($\leq$25drop vs. $\leq$5for rain), potentially as a consequence of the increased object opacity for snow particles. Still, the background error ($\geq$ 75% contribution to corruption robustness) dominates the score, and the robustness ranking for optical flow methods remains stable, whether snow pixels are included in the score calculation or not. Hence, the additional evaluation on snow further substantiates the stability and expressiveness of corruption robustness as an evaluation metric.

### v2

@@ A.2 Links and Checkpoints for Evaluated Models
SEA-RAFT https://github.com/princeton-vl/SEA-RAFT Spring (M)
@@
Ressources and Run Times. We conducted all evaluations on a single NVIDIA RTX A6000 (48 GB) GPU. The evaluation time of models on the RobustSpring data strongly depends on the computational efficiency and requirements of the original models. As a representative example, the optical flow evaluation with RAFT (Teed and Deng, 2020) took 20 hours on the full RobustSpring data.
Once the methods are evaluated and the results uploaded to our benchmark server prototype, the robustness evaluation for an optical flow method takes 13 minutes on our evaluation machine (Intel(R) Xeon(R) Gold 6130 CPU @ 2.10GHz, 8 vCPUs). The robustness evaluation for scene flow takes about 26 minutes, and 7 minutes for stereo.
A.3 RobustSpring Image Corruptions
Below, we provide supplementary information on the image corruptions for the RobustSpring dataset. Besides visualizing further benchmark samples and supplying a video that showcases the space- and time-integration of our corruptions, we also give details on their implementation with a focus on RobustSpring-specific consistencies.
A.3.1 Additional Corruption Benchmark Samples
To complement the benchmark samples in Fig. 3, we show benchmark results on additional corruptions in Fig. 6.
@@ A.3.1 Additional Corruption Benchmark Samples
Figure 6: RobustSpring example frames, complementing Fig. 3. The first row shows clean and corrupted images. The second row shows the left and right disparity maps predicted with LEA Stereo (Cheng et al., 2020). The third row shows the target disparities for forward left, backward left, forward right, and backward right directions from M-FUSE (Mehl et al., 2023a). The fourth row shows optical flow estimates for forward left, backward left, forward right, and backward right from RAFT (Teed and Deng, 2020). All disparities and flows are computed on the corrupted dataset.
A.3.2 Video of Corruptions
With the supplementary material, we include a video that visualizes all corruptions applied to the Spring test dataset.
A.3.3 Corruption Implementation
@@
Defocus Blur. The defocus blur convolves the image with a circular mean filter $C^{\mathit{mean}}$
$\hat{I}=I*C^{\mathit{mean}}_{r},$ (7)
@@
Gaussian Blur. Gaussian blur convolves the image with a Gaussian $C^{\mathit{gauss}}$
$\hat{I}=I*C^{\mathit{Gauss}}_{\sigma},$ (8)
@@
Pixelation. This is achieved by downsampling the image to size $s$, a fraction of its original size, with a box filter $\mathit{box}(I,s)$, followed by upsampling $\mathit{up}(I,s)$ to the size $s$, which is the original size:
$\hat{I}=\mathit{up}(\mathit{box}(I,I\cdot c),I).$ (16)
@@
Snow and Rain. The implementation for snow and rain is based on Schmalfuss et al. (2023), with methodological and performance improvements. On the methodological side, we replaced additive blending with order-independent alpha blending (Meshkin’s Method,  McGuire and Bavoil (2013)) and included global illumination (Halder et al., 2019) in the color rendering. Also, we expanded the monocular two-step motion simulation to multi-step stereo images. On the performance side, we introduce an efficient parallel particle initialization and improve the parallel processing performance.
Fog. Fog is modelled using the Koschmieder model from Wiesemann and Jiang (2016) as
@@
Figure 7: Additional results on accumulated corruption robustness $R^{c}_{\text{EPE}}$ for optical flow models over corruption classes color, blur, quality, and weather. More results are in Fig. 4.
A.4 Additional Experimental Results
@@
A.4.1 Corruption Robustness by Corruption Group
Figure 7 shows the corruption robustness $R^{c}_{\text{EPE}}$ for each optical flow method across the remaining four corruption groups in addition to Fig. 4 in the main paper. It underlines the varying degrees of robustness of the evaluated methods against specific types of corruption.
Figure 8: Accuracy vs. robustness of optical flow methods, measured as EPE and median $R^{c}_{\text{EPE}}$. Small values indicate accurate and robust methods. The dashed line represents a quadratic polynomial fit. Fig. 4(c) shows the average $R^{c}_{\text{EPE}}$.
A.4.2 Accuracy vs. Median Corruption Robustness
In Fig. 8 we show the accuracy-robustness evaluation with the Median corruption robustness, to complement Fig. 4(c) which uses the average corruption robustness. Even though the robustness ranking of methods varies between average and media corruption robustness, cf. Sec. 4.2 for a discussion on ranking differences, the general trend that corruption robustness and accuracy are weakly correlated remains. However, there is still no clear winner, and an accuracy-robustness tradeoff persists among particularly accurate or robust methods.
A.4.3 Schulze Pairwise Comparison Matrix
Table 5: Pairwise comparison matrix for the Schulze method.
SEA-RAFT
@@ A.4.3 Schulze Pairwise Comparison Matrix
SEA-RAFT 0 14 14 14 14 17 17 15 19
GMFlow 6 0 9 14 9 10 16 9 14
MS-RAFT+ 6 9 0 15 11 11 19 12 15
FlowFormer 6 6 5 0 3 12 16 8 13
GMA 6 11 9 17 0 12 20 10 15
SPyNet 3 10 9 8 8 0 13 4 13
RAFT 3 4 1 4 0 7 0 4 9
FlowNet2 5 10 8 12 10 16 16 0 18
PWCNet 1 6 5 7 5 7 11 2 0
The Schulze method is a ranking algorithm used to determine the most preferred candidate based on pairwise comparisons. We include a pairwise comparison matrix in Table 5 for our ranking. The table shows how often the method in row i is better than the method in column j, based on the number of corruptions where method i achieves a lower error than method j. The ranking process consists of the following steps:
A.4.4 Corruption Robustness on Snow
@@
Figure 9: Stability of corruption robustness $R^{c}_{\text{EPE}}$ for snow corruption. Analogous rain results in Fig. 5(a).
@@
Finally, we complement the evaluation of our corruption robustness metric in the presence of rain in Fig. 5(a) with the corresponding evaluation in the presence of snow in Fig. 9. The results with rain translate to snow with the following minor differences: Because snow has less motion blur than rain, it covers fewer pixels (60% of all pixels vs. 90% for rain). For snow, the score drops a bit more than for rain when object pixels are excluded ($\leq$25% drop vs. $\leq$5% for rain), potentially as a consequence of the increased object opacity for snow particles. Still, the background error ($\geq$ 75% contribution to corruption robustness) dominates the score, and the robustness ranking for optical flow methods remains stable, whether snow pixels are included in the score calculation or not. Hence, the additional evaluation on snow further substantiates the stability and expressiveness of corruption robustness as an evaluation metric.
A.4.5 Reproducibility
We compute the noise in the KITTI (Geiger et al., 2012) frames according to Immerkaer (1996), and select the top 10% noisiest frames. The final frames and their noise levels are listed in Tab. 6.
Table 6: Top 10% noisiest KITTI frames with noise estimation from Immerkaer (1996). Frames from kitti15/dataset/training/image_2.
Frame
000061_10.png
000060_10.png
000065_10.png
000068_10.png
000143_10.png
000069_10.png
000174_10.png
000142_10.png
000107_10.png
000175_10.png
000064_10.png
000063_10.png
000066_10.png
000067_10.png
000062_10.png
000055_10.png
000154_10.png
000158_10.png
000157_10.png
000054_10.png
Noise 3.54 3.22 3.07 2.92 2.80 2.79 2.77 2.74 2.69 2.66 2.62 2.61 2.61 2.60 2.48 2.47 2.46 2.46 2.43 2.41
A.5 Depth and Extrinsics Estimation for Corruptions
Disparity and Depth. For depth-dependent corruptions, we estimate disparity maps with MS-RAFT+ (Jahedi et al., 2022; 2024). On the Spring validation set, we find a mean EPE of $1.09$ px and D1-all of $5.60\%$. Mean EPE measures the average Euclidean distance between predicted and ground-truth disparities, while D1-all reports the percentage of pixels with disparity error larger than 3 px or 5% of the ground-truth disparity. These errors are sufficiently small to support realistic augmentations. Note that depth estimates are used by only 3 of the 20 corruption types (fog, snow, and rain; motion blur uses depth only implicitly).
Extrinsics. We estimate camera extrinsics using COLMAP 3.8. As trajectories are not expressed in a common reference frame, quantitative pose metrics such as absolute trajectory error (ATE) or relative pose error (RPE) cannot be meaningfully computed without a rigid alignment step, which itself introduces error. Extrinsics are required for only 2 corruptions (snow, rain). We therefore evaluate them qualitatively: the resulting augmentations exhibit realistic motion behavior for both effects.
A.6 User Study on SSIM Thresholds
The SSIM thresholds of 0.7 for color, blur, quality, and weather corruptions and 0.2 for noise were chosen empirically to yield comparable perceptual corruption strength across categories.
We conducted a focused perceptual study to validate the choice of corruption strengths and their associated SSIM values in RobustSpring. For five representative corruptions (one per main corruption family: contrast, fog, Gaussian blur, Gaussian noise, and pixelation), participants repeatedly viewed a clean Spring frame alongside three corrupted versions. See Figs. 10 and 11 for examples. They then selected the corrupted image that looked most like a “reasonably” corrupted version of the clean frame. This means the image is clearly degraded, inducing a distribution shift, yet still recognizably close to the original. Across 14 user studies, with 10 samples each, for a total of 140 comparisons, we observed agreement rates ranging from 70% to 100% from each user for the SSIM values used in our study. The average agreement was 87.14%, which suggests that the SSIM ranges used for these representative corruptions align with human perception of realistic, moderate corruption strengths.
In the user study, the web interface displayed one clean reference image from Spring and three corresponding corrupted versions of the same image on each page. These images were obtained by applying a single corruption type at three different intensities. We applied thresholds for three levels of severity: low, medium, and high, with corresponding SSIM values of 0.95, 0.7, and 0.2, respectively. The three corrupted images were randomly ordered, and participants were asked to select one image per page that they perceived as the most reasonably corrupted version of the clean image. The corruption should be clearly visible and sufficient to cause a distribution shift but not so strong that the scene appears implausible or too different from the original image. This setup was repeated for the five corruption types listed above, and for each page and participant, the selected intensity was recorded. This allowed us to compute per-participant agreement with the nominal level and the aggregate agreement statistics reported above. In each scenario, the nominal value corresponds to the SSIM value chosen for the respective corruption in our work.
(a) Clean image.
(b) Low severity.
(c) Medium severity (best).
(d) High severity.
Figure 10: Clean image and three corrupted versions with different severity levels for fog used in the user study on SSIM thresholds. “Best” is the one we choose in RobustSpring.
(a) Clean image.
(b) Low severity.
(c) Medium severity.
(d) High severity (best).
Figure 11: Clean image and three corrupted versions with different severity levels for Gaussian noise used in the user study on SSIM thresholds. “Best” is the one we choose in RobustSpring.
A.7 Motion Range of Spring
As discussed in the Spring paper (Mehl et al., 2023b), the dataset covers a wider range of motion than standard benchmarks. While Sintel (Butler et al., 2012) and KITTI (Menze and Geiger, 2015) align with Spring for smaller motions (u: 500-500 to $500$, v: 250-250 to $250$), Spring extends the range up to $1700$ (u) and 750-750 (v). This corresponds to motion magnitudes about 1.5–3 times larger than those in Sintel and KITTI. We highlight this here as one reason for selecting Spring as the foundation of RobustSpring.
A.8 Integration into the Spring Benchmark Website
To illustrate how RobustSpring extends the existing Spring benchmark infrastructure, we provide mock-ups of the website with our additional robustness results. These examples demonstrate how robustness values are integrated alongside existing accuracy metrics.
Overview Page. Fig. 12 shows the modified overview page for optical flow methods. Three new columns report the average robustness scores. This extension allows users to compare methods across both accuracy and robustness at a glance.
Method Detail Page. Fig. 13 displays the detail page for a single optical flow method. In addition to the accuracy results already present in Spring, robustness values are provided per corruption type, as well as aggregated average and median robustness values. This detailed view enables users to assess strengths and weaknesses of individual methods with respect to specific corruptions.
Figure 12: Modified Spring benchmark overview page for optical flow. Newly added robustness columns are highlighted in red.
Figure 13: Modified Spring benchmark method detail page for optical flow. Robustness values are reported for each corruption, as well as aggregated average and median scores. Newly added sections are highlighted in red.